# Deep learning for precipitation nowcasting: A survey from the perspective of time series forecasting

### v1

@@ 4.4 Sliding window
The sliding window approach is a widely used term in time series forecasting tasks. The term refers to a technique where a fixed-size window of consecutive time steps is used as input to predict the next time step or a sequence of future time steps. The window size is an important hyperparameter that needs to be tuned in the forecasting task, which is typically set between 20 min (Ravuri et al., 2021) to 2 hours (h) (Larvor et al., 2020) in nowcasting models. Building on previous studies, DL models can effectively capture temporal dependencies and patterns within precipitation data by adopting the sliding window with intervals exceeding 20 min and considering a fixed number of previous time steps as input to forecast future rainfall intensity.
@@ 7 Recursive strategy
The recursive strategy is commonly used for predicting hazardous weather over a very short term and has been studied with a focus on predicting spatiotemporal motion since leading to biased and suboptimal forecasts, especially for longer forecast lead time. However, the advantage of the recursive forecasting strategy is that it can effectively capture temporal dependencies in the time series, which is commonly composed of recurrent neural networks (Shi et al., 2015; Wang et al., 2017). The strategy can be classified into two types:
@@ 8 Multiple strategy
Advantages and disadvantages. (+) robust against such accumulating errors, (-) but leads to difficulties in capturing temporal dependencies between each frame.
@@
    UNet-based methods: (+) achieve high-performance for multivariate forecasting, (-) but struggle with learning temporal dependencies.
@@ 8.2.1 Sharpness
(a) Schematic of the UNet-based NowcastNet (Zhang et al., 2023). (Left) NowcastNet overview based on UNet (Encoder and Decoder) structure. (Right) A temporal discriminator using multiple convolution kernels. The architecture focuses on a UNet-based model designed to learn space-time dependencies. This is achieved through an evolution network that enables the model to learn both motion flow and intensity change. By combining these sub-networks, the model can effectively capture the dynamic changes in the data over time and space.
@@
(c) Schematic of the Transformer-based MIMO-VP (Ning et al., 2023). (Left) MIMO-VP overview for predicting future time steps. (Right) Encoder-Decoder structure. The architecture combines convolution with time-wise attention to learn space-time dependencies. By leveraging the self-attention mechanism with convolution layers, the model can understand and predict space-time relationships, enabling it to capture complex patterns and dynamics in temporal data. This framework also provides robustness to long-range dependencies, allowing the model to effectively capture complex correlations and patterns.
@@ 8.3.1 Long-term dependency
Transformer-based nowcasting models have achieved remarkable success, demonstrating their ability to effectively capture long-range spatiotemporal dependencies in precipitation nowcasting. Transformer efficiently robust long-range dependencies by tokenizing data into small patches and learning weights for each patch using self-attention. The self-attention mechanism in the Transformer, while powerful for capturing long-range dependencies, may struggle to effectively model short-term local patterns and variations in the data (Lin et al., 2022). Rainformer (Bai et al., 2022) and MS-RadarFormer (Geng et al., 2024) applied a Swin Transformer by tokenizing time series frames to learn both global and local features. They introduced a fully convolutional network with an attention module that effectively extracts local features from rainfall information, which aids in low-to-medium rainfall intensity prediction. However, precipitation data is to fluid masses spreading in various directions, unlike standard videos which often depict moving objects against a stationary background (Fovell and Tan, 1998). To overcome this drawback, MIMO-VP (Ning et al., 2023) introduced a spatiotemporal block aimed at learning high-order relationships between time step queries and their corresponding output sequences. As described in Figure 6(c), the approach involves obtaining sequence-level feature maps using a placeholder embedding method to capture global dependency within the frames. The proposed solution effectively addresses the challenge of global space-time dependency and error accumulation by adopting temporal attention modules with convolution layers. To tackle the permutation-invariant issue inherent to transformers, the network applies temporal positional encoding to input sequences, ensuring smooth and effective processing.
@@ 8.3.3 Periodic component
As scientific data substantiates the importance of Fourier analysis, it is a well-known fact that the decomposition of complex weather signals into their constituent frequency components can better identify and represent important meteorological phenomena (Amon, 1993). The Fourier series representation within the spectral elements enables the efficient handling of periodic or quasi-periodic phenomena, such as those encountered in transitional and turbulent flows (Fournier et al., 2005). The Fourier-based precipitation nowcasting, such as fast Fourier transform (FFT), can help extract relevant features and improve the performance of forecasting models (Wu et al., 2023, 2024). The Fourier representation allows the model to capture complex periodic behaviors through the combination of sine and cosine terms. Passing input points through a Fourier feature mapping enables networks to learn high-frequency functions in a low-dimensional setting. The FFT module is utilized to capture high-dimensional nonlinear physical features, enabling the capture of intricate dynamics within systems without relying on differential equation-based nonlinear features (Wu et al., 2023). Unlike discrete static frames, the FFT approach transforms data from the continuous time domain to the frequency domain, thereby better preserving long-term dependencies in spatiotemporal data (Wu et al., 2024). Fourier transforms have gained prominence in the field of precipitation nowcasting, providing a method to discern intricate physical patterns and capture system dynamics without the need for differential equation-based nonlinear features.
@@
A SAC-LSTM (She et al., 2023) [link] MSE, CE CIKM 101 15 10 6 X CSI, HSS, MSE
@@ 10.4 Standard evaluation protocols
With the increasing number of time series precipitation forecasting models, there is a growing need for a standardized baseline against which different algorithms can be uniformly evaluated. Standardizing the evaluation process is essential to ensure fair and consistent comparisons between different algorithms. In this regard, what should the test period, rainfall thresholds and forecast lead time be? How can we compare new models to the cutting-edge models, such as different resolution models? Is there agreement on whether radar data is the best ground truth for precipitation? Perhaps an estimated value by radar is not the true rainfall, despite the radar-obtained precipitation value potentially being the closest to the actual value. By comparing the performance of different models using standardized evaluation metrics, researchers can identify the strengths and weaknesses of each approach more effectively.

### v2

@@ 4.4 Sliding window
The sliding window approach is a widely used term in time series forecasting tasks. The term refers to a technique where a fixed-size window of consecutive time steps is used as input to predict the next time step or a sequence of future time steps. The window size is an important hyperparameter that needs to be tuned in the forecasting task, which is typically set between 20 min (Ravuri et al., 2021) to 2 hours (h) (Larvor et al., 2020) in nowcasting models. Building on previous studies, DL models can effectively capture temporal dependency and patterns within precipitation data by adopting the sliding window with intervals exceeding 20 min and considering a fixed number of previous time steps as input to forecast future rainfall intensity.
@@ 7 Recursive strategy
The recursive strategy is commonly used for predicting hazardous weather over a very short term and has been studied with a focus on predicting spatiotemporal motion since leading to biased and suboptimal forecasts, especially for longer forecast lead time. However, the advantage of the recursive forecasting strategy is that it can effectively capture temporal dependency in the time series, which is commonly composed of recurrent neural networks (Shi et al., 2015; Wang et al., 2017). The strategy can be classified into two types:
@@ 8 Multiple strategy
Advantages and disadvantages. (+) robust against such accumulating errors, (-) but leads to difficulties in capturing temporal dependency between each frame.
@@
    UNet-based methods: (+) achieve high-performance for multivariate forecasting, (-) but struggle with learning temporal dependency.
@@ 8.2.1 Sharpness
(a) Schematic of the UNet-based NowcastNet (Zhang et al., 2023). (Left) NowcastNet overview based on UNet (Encoder and Decoder) structure. (Right) A temporal discriminator using multiple convolution kernels. The architecture focuses on a UNet-based model designed to learn space-time dependency. This is achieved through an evolution network that enables the model to learn both motion flow and intensity change. By combining these sub-networks, the model can effectively capture the dynamic changes in the data over time and space.
@@
(c) Schematic of the Transformer-based MIMO-VP (Ning et al., 2023). (Left) MIMO-VP overview for predicting future time steps. (Right) Encoder-Decoder structure. The architecture combines convolution with time-wise attention to learn space-time dependency. By leveraging the self-attention mechanism with convolution layers, the model can understand and predict space-time relationships, enabling it to capture complex patterns and dynamics in temporal data. This framework also provides robustness to long-range dependency, allowing the model to effectively capture complex correlations and patterns.
@@ 8.3.1 Long-term dependency
Transformer-based nowcasting models have achieved remarkable success, demonstrating their ability to effectively capture long-range spatiotemporal dependency in precipitation nowcasting. Transformer efficiently robust long-range dependency by tokenizing data into small patches and learning weights for each patch using self-attention. The self-attention mechanism in the Transformer, while powerful for capturing long-range dependency, may struggle to effectively model short-term local patterns and variations in the data (Lin et al., 2022). Rainformer (Bai et al., 2022) and MS-RadarFormer (Geng et al., 2024) applied a Swin Transformer by tokenizing time series frames to learn both global and local features. They introduced a fully convolutional network with an attention module that effectively extracts local features from rainfall information, which aids in low-to-medium rainfall intensity prediction. However, precipitation data is to fluid masses spreading in various directions, unlike standard videos which often depict moving objects against a stationary background (Fovell and Tan, 1998). To overcome this drawback, MIMO-VP (Ning et al., 2023) introduced a spatiotemporal block aimed at learning high-order relationships between time step queries and their corresponding output sequences. As described in Figure 6(c), the approach involves obtaining sequence-level feature maps using a placeholder embedding method to capture global dependency within the frames. The proposed solution effectively addresses the challenge of global space-time dependency and error accumulation by adopting temporal attention modules with convolution layers. To tackle the permutation-invariant issue inherent to transformers, the network applies temporal positional encoding to input sequences, ensuring smooth and effective processing.
@@ 8.3.3 Periodic component
As scientific data substantiates the importance of Fourier analysis, it is a well-known fact that the decomposition of complex weather signals into their constituent frequency components can better identify and represent important meteorological phenomena (Amon, 1993). The Fourier series representation within the spectral elements enables the efficient handling of periodic or quasi-periodic phenomena, such as those encountered in transitional and turbulent flows (Fournier et al., 2005). The Fourier-based precipitation nowcasting, such as fast Fourier transform (FFT), can help extract relevant features and improve the performance of forecasting models (Wu et al., 2023, 2024). The Fourier representation allows the model to capture complex periodic behaviors through the combination of sine and cosine terms. Passing input points through a Fourier feature mapping enables networks to learn high-frequency functions in a low-dimensional setting. The FFT module is utilized to capture high-dimensional nonlinear physical features, enabling the capture of intricate dynamics within systems without relying on differential equation-based nonlinear features (Wu et al., 2023). Unlike discrete static frames, the FFT approach transforms data from the continuous time domain to the frequency domain, thereby better preserving long-term dependency in spatiotemporal data (Wu et al., 2024). Fourier transforms have gained prominence in the field of precipitation nowcasting, providing a method to discern intricate physical patterns and capture system dynamics without the need for differential equation-based nonlinear features.
@@
A SAC-LSTM (She et al., 2023) [link] minmax MSE, CE CIKM 128 31.1 M 5 10 6 X CSI, HSS, MSE
@@ 10.4 Standard evaluation protocols
With the increasing number of time series precipitation forecasting models, there is a growing need for a standardized baseline against which different algorithms can be uniformly evaluated. Standardizing the evaluation process is essential to ensure fair and consistent comparisons between different algorithms. In this regard, what should the test period, rainfall thresholds, and forecast lead time be? How can we compare new models to the cutting-edge models, such as different resolution models? Is there agreement on whether radar data is the best ground truth for precipitation? Perhaps an estimated value by radar is not the true rainfall, despite the radar-obtained precipitation value potentially being the closest to the actual value. By comparing the performance of different models using standardized evaluation metrics, researchers can identify the strengths and weaknesses of each approach more effectively.